What diameter of focusing lens would be needed to achieve a numerical aperture of 1.0 with a working distance of 5 mm.
Answer #1

The f-number of a microscope is the working distance of the objective divided by the aperture diameter. It is defined by the equation

f/D = 1/2(NA)

5mm/D = 1/2*1.0

D = 10mm